# SeeThruFinger: See and Grasp Anything with a Multi-Modal Soft Touch

### v3

@@ 2.2 Formalizing Multi-Modal VBTS via a See-Thru-Network
For an image taken by an in-finger vision camera, such as the case shown in Fig. 1A, the resultant image is essentially the same as the general representation reproduced in Fig. 5. The ambient environment is gray in Fig. 5. If the in-finger camera is used for VBTS architecture with an Illuminated Chamber (IC) shown in Fig. 1B, then any th row and th column pixel of the image is regarding the state of the soft membrane captured within the camera’s viewing range, containing zero information about the scene before contact begins.
@@
However, for an image taken by an in-finger camera for VBTS architecture with a See-Thru-Network (STN), such as the case shown in Fig. 1C, the resultant image contains the scene occluded by the Soft Polyhedral Network (SPN). We denote the image as if the SPN is in a Static State before any contact. We can add a hat to use the symbol to indicate a particular case if the scene is empty, captured as uniform pixel features. Here, the subscript timestep is removed to simplify the representation. Since the valuable information captured in this image is purely regarding the SPN, one can add a transparency channel to obtain , represented in white color in Fig. 5. Pixel data of the SPN is preserved in this added data channel, but the pixel data of the empty scene is set to transparent. Such an operation would produce a matrix that can be visualized as a Static Mask of the segmented SPN. Element-wise matrix multiplication of image and this Static Mask will produce masked image . Note that performing the same operation using raw image will produce the same masked image instead. This also explains the SPN’s See-Through feature in the in-finger vision image occlusion process. The above process generally applies to scenarios when the SPN is in a Static State, meaning that no deformation is induced as there is no contact event. One can use classical image processing techniques to obtain the above results.
@@
Using the symbols above, we formalize three research problems of this study for a generic scenario of reactive object grasping where both 1) scene-wise visual perception for object detection and 2) object-centric tactile sensing for force-and-torque sensing are considered.
@@
During contact-based physical interactions, can we infer high-precision tactile sensing of 6D forces and torques in a markerless way?
Figure 6 illustrates these three research problems when applying the soft polyhedral networks with in-finger vision as the fingertips. The following section will provide further details on addressing these three research problems by leveraging state-of-the-art algorithms from computer vision. Section 4 offers a full experimental demonstration of reactive grasping using the SeeThruFingers only. All three research problems are addressed using the See-Thru-Network architecture without external cameras or dedicated force-and-torque sensors. Section 5 discusses the further expansion of these three research problems into a multi-modal approach for robotics using vision-based tactile sensing via SeeThruFingers.
3 Methods & Results
3.1 Markerless, Real-time, Large-Scale Deformation Tracking
@@
The first problem is to enable markerless real-time segmentation tracking of the SPN’s adaptive deformation during contact, which always covers a considerable amount of the image area with large-scale spatial deformations. This problem is equivalent to video object segmentation in computer vision. A significant challenge of this problem is to efficiently manage the feature memory model to avoid memory explosion with minimized performance decay for long-term predictions, where the latest development in a Unified Memory architecture, the XMem model, exhibits state-of-the-art performances on long-video benchmarking datasets [5].
@@
In this study, we experimented with the XMem model for the in-finger video stream, as shown in Fig. 7, achieving a real-time implementation (30 Hz) of the SPN’s large-scale deformation mask in a markerless way. This is realized by using a Static Mask of the SPN as the initial tracking target and the in-finger vision image of the SPN after contact. Before any contact event, this Static Mask remains the same and can be conveniently obtained manually or using advanced segmentation algorithms such as SAM [4].
@@
We evaluated this mask segmentation over a test dataset of 1,000 samples with the ground truth masks. The region (J) and boundary (F) measures [39] are 0.975 and 0.997, respectively, showing excellent reliability in tracking the soft finger’s spatial deformation masks in a markerless way. In applications, we found that the XMem’s mask tracking of the SPN’s deformation is not limited to the Dynamic States but also applicable to the Static States of the SPN. We successfully implemented markerless and real-time deformation tracking of the SPN whenever the SeeThruFinger architecture is enabled, alleviating the need for artificial markers reported in previous VBTS literature. In other words, this process can be alternatively used to detect the occurrence of a contact event. This leads to the formal representation of this implementation as the following, addressing the first research problem in Fig. 6.
3.2 In-Finger Visual Perception from Scene Inpainting
@@
The second problem is reconstructing the scene using occluded images captured by the in-finger vision for visual perception, such as object detection, providing the equivalent capability to an external “Eye” camera. Due to the SPN’s See-Through feature, this problem becomes comparable to the classical inpainting problem, a conservation technique to fill in damaged, deteriorated, or missing parts of an artwork to present a complete and coherent image. Video inpainting has recently attracted significant interest from the computer vision research community, where learning-based methods exhibit superior performances in filling in missing regions of a video with spatially and temporally coherent contents [40].
@@
Therefore, we propose to move around the SeeThruFinger shown in Fig. 8A to obtain redundant information about the scene by scanning sequential in-finger images from different angles about the same scene, accumulating sufficient visual features from the See-Through pixels to inpaint the occluded scene. Figure 8B illustrates a hard-coded head-up movement of the gripper for the in-finger camera to capture a short video in which its viewing range sweeps across the tabletop scene. The predefined movement can also be set to linear movement above the robot’s working area with a fixed orientation. The first row of Fig. 8C shows the in-finger vision of the scene without attaching the SPN on the SeeThruFingers, the second row with the soft finger occluding the scene, and the third row for the corresponding frames of the inpainted scene. Only the right half of the image with objects is inpainted for efficient processing. We ensured that all objects have been captured in at least one or more frames in the video, providing context for reconstructing frames where SPN blocks objects.
@@
Then, we implemented the End-to-End framework for the Flow-Guided Video Inpainting (E2FGVI) algorithm [41], a state-of-the-art algorithm for video inpainting. To facilitate the inpainting process, we obtain a Static Mask of the SPN in Fig. 8D using threshold segmentation against a clean background, denoted as . Results show that the SPN occludes about 53.6% of the in-finger image pixels. Alternatively, we can interactively obtain the same mask using the Segment Anything Model (SAM) by clicking for comparison in an actual scene [4]. Visual perception is usually performed before the contact event when for an object manipulation pipeline. During this period, the SPN is in a Static State, meaning that the resultant mask remains unchanged as the initial mask . The E2FGVI algorithm has been reported with a high Structural Similarity Index (SSI) measure of 0.97 on two public datasets, YouTube-VOS and DAVIS, under stationary masks.
Figure 8E shows the scene inpainting pipeline to implement the E2FGVI algorithm using in-finger video clip and the SPN’s Static Mask to generate inpainted scene with objects on the tabletop. This inpainting process is represented as
The inpainted images managed to reconstruct sufficient details of the occluded objects and the scene, as shown in the dashed white boxes in Fig. 8C.
We further implemented a variety of visual perception tasks based on the inpainted scene to demonstrate the feasibility of in-finger visual perception. Firstly, we use the Real-Time DEtection TRansformer (RT-DETR), a state-of-the-art object detection algorithm [42]. This object detection process is represented as
@@
The example shown in Fig. 8F verifies the feasibility of visual perception using an inpainted scene from in-finger vision occluded by the SPN, producing object classification and bounding box center that can be later used for motion planning. Equations (2) and (3) are the formal representation of this implementation, addressing the second research problem in Fig. 6.
3.3 Markerless Tactile Sensing via Deformation Mask
@@
The third problem is to infer high-performing tactile sensing in a markerless way using in-finger vision of the SPN. This representation problem is usually resolved by adding artificial markers to the soft medium during physical interactions, which VBTS architectures share using either an Illuminated Chamber or a See-Thru-Network. Previous research also explored fiducial markers inside the SPN, achieving high-performing tactile sensing, especially when considering the SPN’s static and dynamic viscoelasticity [33].
@@
This study uses the SPN masks in the Dynamic States obtained from segmented tracking as the model input to infer tactile sensing in 6D forces and torques. We adopted the Supervised Variational AutoEncoder (SVAE) for this task, and Fig. 9A shows the experimental setup for collecting true labels of 6D forces and torques at the finger base using nano25 by ATI for calibration. The in-finger camera and FT sensor were connected to a laptop for data streaming. We collected 40,000 synchronized pairs of images and FT readings while manually compressing the two SPN fingers from various directions and contact locations. We used an 8:2 train-valid split and trained the SVAE model for 50 epochs. To avoid excessive good validation scores due to the potential sample dependence, we separately collected a test dataset of 1,000 samples.
@@
Figure 9B illustrates the SVAE model. The SPN masks segmented in Dynamic States and the mask template are fed to the ResNet18 encoder and decoder module to reconstruct the in-finger image . The latent space is set to 32-dimension. At the same time, we design an auxiliary supervised regression task of the 6D forces and torques simultaneously from the learned latent geometry representation. This leads to the formal representation of this implementation as the following, addressing the third research problem in Fig. 6.
The overall loss consists of reconstruction loss, the Mean-Squared Errors (MSE) loss of force and torque prediction, and the Kulback-Leibler divergence. The average time for a single inference is 1.9 ms on an NVidia 3080 Ti Laptop GPU. The Mean Absolute Errors of the validation and test dataset are reported in Table 1, indicating good generalization of the learned model.
@@
| Validation dataset | 0.241 | 0.212 | 0.14 | 0.024 | 0.027 | 0.005 |
| Test dataset | 0.328 | 0.283 | 0.24 | 0.027 | 0.036 | 0.005 |
@@
Figure 10A compares the test dataset’s prediction and ground truth in time series. Figure 10B shows the scatter plot of such comparison in the validation dataset with excellent scores over 0.985 in all components except . The relatively weak perforce in is due to the limited adaptability along the -axis. We also evaluate the resultant force in the plane and find the mean magnitude and orientation error are 0.286 N and 2.6 degrees, respectively.
4 Reactive Grasping with SeeThruFingers
@@
This section demonstrates the application of the proposed SeeThruFinger architecture, shown in Fig. 3, in an object-picking task, shown in Fig. 4, using VBTS with a See-Thru-Network to achieve object detection and tactile sensing for reactive grasping simultaneously.
@@
The task begins with visual perception of the scene using the SeeThruFinger for object classification and localization. During the first attempt, the gripper will attempt to pick up the object at the center of the bounding box and a fixed rotation angle and depth perpendicular to the table. We intentionally used a simpler algorithm without the optimal orientation to test reactive grasping with tactile sensing using the SeeThruFinger. In our previous work [43], it was found that parallel finger configurations lead to poorer grasp performance even with soft adaptation when the grasp angle is not optimized for target objects, especially those with large aspect ratios such as pringles can and spatula in the YCB dataset [38]. Rearranging the soft fingers in a radial configuration leads to a more robust grasp [43], and a significant portion of existing commercial grippers are designed with two fingers in parallel. Hence, it is worthwhile to incorporate tactile sensing for a more robust parallel grasp. The objects used for the demonstration are aspect ratios between 2 to 4 and various weights.
@@
While approaching the initial grasping pose, the SeeThruFinger detects if it has contact with the object by inferring real-time tactile feedback in 6D forces and torques based on the finger’s whole-body deformation. Due to the network design, the finger can deform twistedly to generate torque about the -axis. When is detected to be larger than 0.05 Nm, the object orientation is severely different from the gripper’s initial pose, suggesting a less ideal grasping that can be optimized through regrasping by turning the wrist joint reversely to reduce for an enhanced adaptation and grasping robustness. When is within the threshold, the gripper closes its fingers until the gripping force reaches the predefined maximum gripping force of 3 N. Then, the gripper lifts the object and drops it into the placing box. We repeated the process until all the five objects were cleared from the table.
@@
Figure 11A shows the recorded 6D FT sensing data using the SeeThruFinger for the table-cleaning task, where the six peaks in indicate the grasping of the six test objects. The inpainted scene using SeeThruFinger is shown in Fig. 11B, where the occluded regions are inpainted with coherent details, and the objects’ overall contours are well reconstructed. Although some surface textures generated are blurry, they are sufficiently accurate to reflect their actual textures. Figure 11C shows object detection results based on the inpainted scene, showing bounding boxes for each detected object. The most exciting result is the regrasping of the Bosche electrical screwdriver shown in Fig. 11D. By inspecting the detailed FT history in Fig. 11A, we identified two regrasping attempts of the Bosche electrical screwdriver performed by the SeeThruFinger based on tactile sensing and adaptive grasping.
@@
Figure 11E is an enlarged view of the region marked in a red dashed box, as shown in Fig. 11A. When the SeeThruFinge performed the first grasping attempt in Fig. 11E(i), the finger detected the greater than 0.05 Nm in the negative direction. Then, in Fig. 11E(ii), the first reactive grasping begins rotating the gripper about the -axis to reduce . During the process, the screwdriver suddenly turns to another side and temporarily loses contact with the finger in Fig. 11E(iii), as both and dropped to 0 suddenly. In Fig. 11E(iv), the gripper keeps closing, attempting to detect contact with the screwdriver again to keep increasing for an enhanced grasping pose. Next, the SeeTruFinger detected contact with the screwdriver again in Fig. 11E(v), marking the starting point of the second reactive grasping attempt. The SeeThruFinger repeated the same procedure by performing the second reactive grasping in Fig. 11E(vi), rotating the gripper about the -axis until the detected is reduced towards 0. As a result, the screwdriver is secured between the fingers and moved toward the drop box in Fig. 11E(vii). Finally, the SeeThruFinger dropped the screwdriver with zero values detected from tactile sensing using in-finger vision, as shown in Fig. 11E(viii).
5 Towards a Multi-Modal Soft Touch
This section discusses the intuitive understanding of the three research problems illustrated in Fig. 6 regarding the SeeThruFinger (Fig. 1C) and the See-Thru-Network architecture (Fig. 3) towards multi-modal perception with a soft touch.
5.1 Markerless, Real-Time & Large-Scale Deformation Tracking of the See-Thru-Network

### v4

@@ 3.2 Formalizing Multi-Modal VBSeeThruP
Suppose the in-finger camera is used for VBDeformP with an Illuminated Chamber (IC), the pixels of the image are related to the state of the soft membrane captured within the camera’s viewing range, containing zero information about the scene before contact begins [35, 21].
@@
However, for an image taken by an in-finger camera for VBSeeThruP with a See-Thru-Network (STN), the resultant image contains the scene occluded by the Soft Polyhedral Network (SPN). We denote the image as if the SPN is in a Static State before any contact. We can add a hat to use the symbol to indicate a particular case if the scene is empty, captured as uniform pixel features. Here, the subscript timestep is removed to simplify the representation. Since the valuable information captured in this image is purely regarding the SPN, one can add a transparency channel to obtain , represented in white color in Fig. 3. Pixel data of the SPN is preserved in this added data channel, but the pixel data of the empty scene is set to transparent. Such an operation would produce a matrix that can be visualized as a Static Mask of the segmented SPN. Element-wise matrix multiplication of image and this Static Mask will produce masked image . Note that performing the same operation using raw image will produce the same masked image instead. This also explains the SPN’s See-Through feature in the in-finger vision image occlusion process. The above process generally applies to scenarios when the SPN is in a Static State, meaning that no deformation is induced, as there is no contact event. One can use classical image processing techniques to obtain the above results.
@@
Using the symbols above, we formalize three research problems of this study for a generic scenario of reactive object grasping where both 1) scene-wise visual perception for object detection and 2) object-centric force/torque sensing are considered, as shown in Fig. 4 and Supplementary Video S1.
@@
During contact-based physical interactions, can we infer high-precision 6D forces and torques in a markerless way?
3.2.1 Markerless, Real-time, Large-Scale Deformation Tracking
@@
The first problem is to enable markerless real-time segmentation tracking of the STN’s adaptive deformation during contact, which always covers a considerable image area with large-scale spatial deformations. This problem is equivalent to video object segmentation in computer vision. A significant challenge in addressing this problem is to efficiently manage the feature memory model, avoiding memory explosion while minimizing performance decay for long-term predictions. The latest development in a Unified Memory architecture, the XMem model, exhibits state-of-the-art performance on long-video benchmarking datasets [65].
@@
In this study, we experimented with the XMem model for the in-finger video stream, as shown in Fig. 5, achieving a real-time implementation (30 Hz) of the STN’s large-scale deformation mask in a markerless way.
This is realized by using a Static Mask of the STN as the initial tracking target and the in-finger vision image of the STN after contact. Before any contact event, this Static Mask remains the same and can be conveniently obtained manually or using advanced segmentation algorithms such as SAM [19]. The following formal representation addresses the first research problem in Fig. 4.
@@
We evaluated this mask segmentation over a test dataset of 1,000 samples, using the ground truth masks as a reference. The region (J) and boundary (F) measures [66] are 0.975 and 0.997, respectively, showing excellent reliability in tracking the soft finger’s spatial deformation masks in a markerless way. In applications, we found that XMem’s mask tracking of the STN’s deformation is not limited to the Dynamic States but also applicable to the Static States of the STN. We successfully implemented markerless and real-time deformation tracking of the STN whenever our proposed architecture is enabled, alleviating the need for artificial markers reported in previous VBDeformP literature. This process can also be used to detect the occurrence of a contact event, which will be further elaborated upon in the Results section.
3.2.2 In-Finger Visual Perception from Scene Inpainting
@@
The second problem is reconstructing the scene using occluded images captured by the in-finger vision for visual perception, such as object detection, providing the equivalent capability to an external “Eye” camera. Due to the SPN’s See-Through feature, this problem becomes comparable to inpainting. Video inpainting has recently attracted significant research interest, where learning-based methods exhibit superior performances in filling in missing regions of a video with spatially and temporally coherent contents [67].
@@
We propose moving the soft finger shown in Fig. 6A to obtain redundant information about the scene by scanning sequential in-finger images from different poses of the same scene, accumulating sufficient visual features from the See-Through pixels to inpaint the occluded scene. Figure 6B illustrates a hard-coded head-up movement of the gripper for the in-finger camera to capture a short video in which its viewing range sweeps across the tabletop scene. The predefined movement can be linear above the robot’s working area with a fixed orientation. The first row of Fig. 6C shows the in-finger vision of the scene without attaching the STN, the second row shows the STN occluding the scene, and the third row shows the corresponding frames of the inpainted scene. Only the right half of the image with objects is inpainted for efficient processing. We ensured that all objects have been captured in at least one frame in the video, providing context for reconstructing frames where STN blocks objects.
@@
We adopted the Flow-Guided Video Inpainting (E2FGVI) algorithm [68], a state-of-the-art algorithm for video inpainting. To facilitate the inpainting process, we obtain a Static Mask of the STN in Fig. 6D using threshold segmentation against a clean background, denoted as . Results show that the STN occludes about 53.6% of the in-finger image pixels. Alternatively, we can interactively obtain the same mask using the Segment Anything Model (SAM) by clicking for comparison in an actual scene [19]. Visual perception is usually performed before the contact event when for an object manipulation pipeline. During this period, the STN is in a Static State, meaning that the mask remains unchanged as the initial mask . The E2FGVI algorithm has been reported with a high Structural Similarity Index (SSI) measure of 0.97 on two public datasets, YouTube-VOS and DAVIS, under stationary masks. Figure 6E shows the scene inpainting pipeline to implement the E2FGVI algorithm using in-finger video clip and the STN’s Static Mask to generate inpainted scene with objects on the tabletop. This inpainting process is represented as
The following focuses on implementing the reactive grasping task in Fig. 1. We use the Real-Time DEtection TRansformer (RT-DETR), a state-of-the-art object detection algorithm [69]. This object detection process is represented as
@@
The example shown in Fig. 6F verifies the feasibility of visual perception using an inpainted scene from in-finger vision occluded by the SPN, producing object classification and bounding box center that can be later used for grasp planning. Equations (2) and (3) are the formal representation of this implementation, addressing the second research problem in Fig. 4.
3.2.3 Markerless Contact Perception via a Deformation Mask
@@
The third problem is to infer high-performing contact-based perception markerlessly using in-finger vision. This representation problem is usually resolved by adding artificial markers to the soft medium, which VBSeeThruP shares using either an Illuminated Chamber or a See-Thru-Network. Previous research has also explored fiducial markers within the STN, achieving high-performing proprioceptive perception, particularly when considering the STN’s static and dynamic viscoelastic properties [18].
@@
This study utilizes the STN masks in the Dynamic States, obtained from segmented tracking, as the model input to infer 6D forces and torques. We adopted the Supervised Variational AutoEncoder (SVAE) for this task, and Fig. 7A shows the experimental setup for collecting true labels of 6D forces and torques at the finger base using nano25 by ATI for calibration. The in-finger camera and FT sensor were connected to a laptop for data streaming. We collected 40,000 synchronized pairs of images and FT readings while manually compressing the two STN fingers from various directions and contact locations. We used an 8:2 train-valid split and trained the SVAE model for 50 epochs. To avoid excessive good validation scores due to the potential sample dependence, we separately collected a test dataset of 1,000 samples.
@@
Figure 7B illustrates the SVAE model. The STN masks segmented in Dynamic States and the mask template are fed to the ResNet18 encoder and decoder module to reconstruct the in-finger image . The latent space is set to 32 dimensions. At the same time, we design an auxiliary supervised regression task of the 6D forces and torques simultaneously from the learned latent geometry representation. This leads to the formal representation of this implementation, addressing the third research problem in Fig. 4.
@@
| Validation | 0.241 | 0.212 | 0.14 | 0.024 | 0.027 | 0.005 |
| Test | 0.328 | 0.283 | 0.24 | 0.027 | 0.036 | 0.005 |
The overall loss comprises reconstruction loss, Mean-Squared Error (MSE) loss for force and torque prediction, and Kullback-Leibler divergence. The average time for a single inference is 1.9 milliseconds on an NVIDIA GeForce RTX 3080 Ti Laptop GPU. The Mean Absolute Errors of the validation and test datasets are reported in Table 2, indicating good generalization of the learned model.
@@
Figure 7C compares the test dataset’s prediction and ground truth in time series. Figure 7D shows the scatter plot of such comparison in the validation dataset with excellent scores over 0.985 in all components except . The relatively weak perforce in is due to the limited adaptability along the -axis. We also evaluate the resultant force in the plane and find that the mean magnitude and orientation error are 0.286 N and 2.6 degrees, respectively.
3.3 Ablation Study
This section presents an ablation study demonstrating how VBSeeThruP contributes to a multi-modal grasping pipeline without requiring additional or external devices. In particular, we investigate how in-finger vision and soft compliance from VBSeeThruP improve the reliability and generalization in a pick-and-place scenario. Please refer to Supplementary Video S2 for further details.
We begin by performing depth detection of table height relative to the robot base frame. Specifically, we tilt the gripper vertically downward until it contacts the table with a force of . Next, we calibrate the hand-eye in the desktop’s 2D plane to obtain the perspective transformation between VBSeeThruP’s in-finger camera space and the robot base frame.
For the calibration process, we place four oranges (or any objects of similar size) on the table and detect their centers in the image plane using YOLO11 [70] (see Fig. 8A). An operator then manually bends the soft finger (for horizontal - movement) to position the gripper over each orange. Once the gripper is centered above an orange, we record the corresponding world coordinates in the robot base frame by twisting the soft finger (clockwise and counter-clockwise to mark the start and end of data recording). The overall setup is shown in Figs. 8B&C, and a full demonstration is provided in the supplementary video. Notably, our method solely relies on VBSeeThruP for depth detection, whereas a typical solution would require an external depth camera.
We employ the following control policy to facilitate intuitive robot teaching:
where denotes the translational velocity in the -th direction, and the corresponding finger deformation. For convenience, we also define two additional commands triggered by the torque :
Using the pairs of recorded images and robot-world coordinates, we compute the homography matrix , enabling transformations between the in-finger vision space and the robot base frame.
3.3.1 Abalation Experiment Setup
We conducted a pick-and-place sorting task with six objects sharing similar geometry but different weights, including 1) Two hollow plastic balls ( radius): one empty and one filled with a metal block; 2) Two plastic capsules ( radius, length): one empty and one containing of jelly balls; 3) One orange ( radius, ); and 4) One carrot ( radius, ).
We tested four different ablation conditions:
1.
SPN Only: We use the SPN’s adaptive grasping with no in-finger vision or FT. The grasp position and rotation are randomly generated, and the gripper’s closing width is fixed at for all objects.
2.
SPN + In-finger Vision: We augment the SPN with soft fingers’ in-finger vision for scene inpainting and YOLO11-based [70] object localization. The grasp width remains fixed at .
3.
SPN + In-finger Vision + Inferred 6D FT: Besides in-finger vision, the system leverages the soft finger’s deformable sensor for estimating 6D forces and torques. We use reactive grasp control to maintain a constant grasp force of () and continuously monitor grasp force during transport. If the force drops to zero (e.g., due to an external disturbance), the system will attempt to regrasp.
4.
Rigid + External Vision + Wrist FT: We replace soft fingers with the original rigid fingertips and rely on an external camera for object localization. The gripper width is set based on the object’s oriented bounding box, and wrist-mounted force/torque sensing on the UR10e provides feedback.
We evaluate each setup on three aspects:
Pick-and-place success: The ability to pick objects from the tabletop and place them into a designated box.
Weight classification before placement: In the third ablation condition, we aggregate the two readings from each soft finger and estimate the object mass as
(7) We use the reading from UR10e’s wrist FT sensing in the fourth ablation condition.
Recovery from human disturbance: We manually remove the carrot from the gripper during the transport (Fig. 8D) to test whether the robot detects the event and regrasps successfully.
3.3.2 Abalation Results and Discussion
Each experiment scenario was repeated three times. We observed that even in the most straightforward configuration (SPN Only), objects could often be picked without damaging them or triggering safety stops, owing to the compliance offered by the soft finger. For example, the soft, deformable fingertips are easily adapted to objects from different angles, as shown in Fig. 8E (grasping an orange from the side with bending and twisting deformations). In contrast, using the original rigid fingertips without carefully specifying object dimensions frequently caused collisions and emergency stops.
When in-finger vision was added (SPN + In-finger Vision), the grasp location and orientation were derived from YOLO11, significantly improving the success rate over random grasps. Nevertheless, the lack of force feedback made the system less robust under disturbance, causing the robot to fail regrasp attempts in several trials (shown in the supplementary video). In one case, the fingertip was misaligned and exerted excessive force on a plastic ball, causing it to slide out of the gripper.
The configuration with SPN + In-finger Vision + Inferred 6D FT overcame these issues using 6D force/torque estimates for closed-loop grasping. As a result, it achieved a 100% success rate under the same disturbance conditions. It also enabled weight classification to differentiate between empty balls and allowed the system to adaptively regrasp objects if an object fell was detected.
Comparatively, the Rigid + External Vision + Wrist FT setup offered an unobstructed camera view of the workspace but lacked the compliance of soft fingers. Small errors in object detection or hand-eye calibration frequently led to inappropriate grasp forces, sometimes resulting in damaged objects (Fig. 8F) or failed grasps. All experiment outcomes are summarized in Fig. 8G.
The results indicate that VBSeeThruP’s in-finger vision and compliance significantly improve the success rate and robustness of pick-and-place tasks over alternative setups. The robot achieves reliable picking, weight classification, and effective regrasping after disturbances without requiring external cameras or specialized sensors. Even when paired with an unobstructed external vision system, rigid fingers are more likely to cause failure or exert excessive force, underscoring the advantage of soft fingers with built-in sensing.
4 Results and Discussion
4.1 Learning Reactive See-Through Grasping
@@
This section demonstrates the application of the proposed architecture for learning reactive see-through grasping, shown in Fig. 1, in an object-picking task, shown in Fig. 2, using VBSeeThruP with a See-Thru-Network to achieve object detection and 6D forces and torques for reactive grasping simultaneously. Please refer to Supplementary Video S3 for a video demonstration.
@@
The task begins with visual perception of the scene using the VBSeeThruP for object classification and localization. During the first attempt, the gripper will attempt to pick up the object at the center of the bounding box and a fixed rotation angle and depth perpendicular to the table. We intentionally used a simpler algorithm without the optimal orientation to test reactive grasping.
@@
While approaching the initial grasping pose, VBSeeThruP detects if it has contact with the object by inferring real-time 6D forces and torques based on the finger’s whole-body deformation. Due to the network design, the finger can deform in a twisted manner to generate torque about the -axis. When is detected to be larger than 0.05 Nm, the object orientation is severely different from the gripper’s initial pose, suggesting a less ideal grasping that can be optimized through regrasping by turning the wrist joint reversely to reduce for an enhanced adaptation and grasping robustness. When is within the threshold, the gripper closes its fingers until the gripping force reaches the predefined maximum gripping force of 3 N. Then, the gripper lifts the object and drops it into the placement box. We repeated the process until all five objects were cleared from the table.
@@
Figure 9A shows the recorded 6D FT sensing data using VBSeeThruP for the table-cleaning task, where the six peaks in indicate the grasping of the six test objects. The inpainted scene using VBSeeThruP is shown in Fig. 9B, where the occluded regions are inpainted with coherent details, and the overall contours of the objects are well reconstructed. Although some surface textures generated are blurry, they are sufficiently accurate to reflect their actual textures. Figure 9C shows object detection results based on the inpainted scene, showing bounding boxes for each detected object. The most exciting result is the regrasping of the Bosche electrical screwdriver shown in Fig. 9D. By inspecting the detailed FT history in Fig. 9A, we identified two regrasping attempts of the Bosche electrical screwdriver performed by the VBSeeThruP with the soft fingers.
@@
Figure 9E is an enlarged view of the red dashed box region in Fig. 9A. When we performed the first grasping attempt in Fig. 9E(i), the finger detected the greater than 0.05 Nm in the negative direction. Then, in Fig. 9E(ii), the first reactive grasping begins rotating the gripper about the -axis to reduce . During the process, the screwdriver suddenly turns to another side and temporarily loses contact with the finger in Fig. 9E(iii), as both and dropped to 0 suddenly. In Fig. 9E(iv), the gripper keeps closing, attempting to detect contact with the screwdriver again to keep increasing for an enhanced grasping pose. Next, the soft finger detected contact with the screwdriver again in Fig. 9E(v), marking the starting point of the second reactive grasping attempt. The soft finger repeated the same procedure by performing the second reactive grasping in Fig. 9E(vi), rotating the gripper about the -axis until the detected is reduced towards 0. As a result, the screwdriver is secured between the fingers and moved toward the drop box in Fig. 9E(vii). Finally, the soft finger dropped the screwdriver with forces and torques approaching zero, as shown in Fig. 9E(viii).
4.2 Towards a Multi-Modal Soft Touch vis VBSeeThruP
Here, we discuss the extended modalities perceivable via our proposed architecture in Fig. 1, providing an intuitive understanding of the three research problems illustrated in Fig. 4 that aim to achieve multi-modal perception with a soft touch. Please refer to the following sections and Supplementary Video S4 for further details.
Using a single sensory input from the in-finger vision featuring See-Thru-Networks, we present a multi-modal approach towards robotic perceptual embodiment via VBSeeThruP with a soft touch, as shown in Fig. 10. Our proposed architecture can be generalized to further generate various outputs for extrinsic perception, including inpainted scenes, object detection, depth sensing, and scene segmentation. For intrinsic perceptions, we demonstrate the implementation of our architecture, which generates outputs including masked deformation tracking, 6D FT sensing in real-time with high accuracy, and contact event detection. Further expansion to other intrinsic or extrinsic modalities can also be implemented.
4.2.1 Markerless, Real-Time & Large-Scale Deformation Tracking